if the ratio of sum of n terms of  two APs is (3n+8):(7n+15), then the ratio of 12th terms is??

Grade:12

2 Answers

vikas askiitian expert
13 years ago

Sn = n/2(2a + (n-1)d)

(Sn)1/(Sn)2 = [2a1 + (n-1)d1] / [2a2 + (n-1)d2]

 

                = [a1 + (n-1)d1/2 ] / [a2 + (n-1)d2/2]

this expression will become ratio of nth term if , (n-1)/2 is made (k-1) ...

(Sn)1/(Sn)2 = 3n+8/7n+15                       (given)

for (n-1)/2 = k-1         or        n = 2k-1

(Sn)1/(Sn)2 = (Tn)1/(Tn)2 = 3(2k-1) + 8/7(2k-1) + 15 = (6k + 5)/(14k + 8)

now put k=12 , this will give ration of 12th term

T12 = 77/176

approve my ans if u like it

Aiswarya Ram Gupta
13 years ago

thnx